OpenStudy (anonymous):

what is the period of y=2cos4x

OpenStudy (anonymous):

period of y=cos(ax) is \[\Large \frac{2\pi}{a}\] here a=4

OpenStudy (anonymous):

so the period is \[\pi\] ?